OpenStudy (is3535):

What is the best order-of-magnitude estimate for 58,100? A. 103 B. 104 C. 105 D. 106

OpenStudy (anonymous):

\(10^3=1000\); \(10^4=10,000\); \(10^5 = 100,000\); \(10^6 = 1,000,000\). Which of these is 58,100 closest to?

OpenStudy (is3535):

C

OpenStudy (is3535):

@ospreytriple

OpenStudy (anonymous):

Correct.

OpenStudy (is3535):

can you help me in one more

OpenStudy (anonymous):

Please post.